A parabola has x-intercepts of -5, and 3 and passes through the point (1,-12). Find its equation.
Answer by stanbon(75887)

A parabola has x-intercepts of -5, and 3 and passes through the point (1,-12). Find its equation.
------
y = a(x+5)(x-3)
---
Since f(1) = -12,
a(1+5)(1-3) = -12
a(6)(-2) = -12
a = 1
-----
Ans:; y = x^2+2x-15
----------------
Cheers,
Stan H.
-------------